Please answer all questions show calculations of how you got the answer this is due in the morning your help is appreciated thank you :)

1.) An inflated helium balloon with a volume of 0.55 L at sea level (1.0 atm) is allowed to rise to a height of 6.5 km, where the pressure is about 0.40 atm. Assuming that the temperature remains constant, what is the final volume of the balloon?

2.) The amount of gas inside the ballon and its temperature remain constant but both the pressure and volume change. What gas law do you need?

3.) Argon is an inert gas used in lightbulbs to retard the vaporization of the tungsten filament. A certain lightbulb containing aragon 1.20 atm and 18 degrees celcius is heated to 85 degree celcius at constant volume. Calculate its final pressure (in atm).

4.) A small bubble rises from the bottom of a lake, where the temperature and pressure are 8°C and 6.4 atm, to the water’s surface, where the temperature is 25°C and the pressure is 1.0 atm. Calculate the final volume (in mL) of the bubble if its initial volume was 2.1 mL.
Answer #1

1)According to ideal gas law PV = nRT

Where P - Pressure , V - volume , n = numkber of moles , R = gas constant , T - Temperature

As the gas is the same n will remain constant , T also kept constant

So PV = constant

or P1V1 = P2V2

Where

P1 = initial pressure = 1.0 atm

V1 = initial volume = 0.55 L

P2 = final pressure = 0.40 atm

V2 = final volume = ?

Plug the values we get V2 = (P1V1) / P2

= 1.375 L

-----------------------------------------------------------------------------------------------------------

(2) The relation PV = constant represents the Boyle's law

------------------------------------------------------------------------------------------------------------

(3) As n , V are kept constant P is proportional to T

OR     P1 / T1 = P2 / T2

Where

P1 = initial pressure = 1.20 atm

T1 = initial temperature = 18 oC = 18 + 273 = 291 K

P2 = final pressure = ?

T2 = final temperaturte = 85 oC = 85+273 = 358 K

Plug the values we get P2 = (P1T2) / T1

                                      = 1.48 atm

------------------------------------------------------------------------------------------------------

(4)The relation is (P1V1) /T1 = ( P2V2) / T2

Where

P1 = initial pressure = 6.4 atm

T1 = initial temperature = 8 oC = 8 + 273 = 281 K

V1 = initial volume = 2.1 mL

P2 = final pressure = 1.0 atm

T2 = final temperaturte = 25 oC = 25+273 = 298 K

V2 = final volume = ?

Plug the values we get

V2 = (P1V1T2)/(P2T1) = 14.25 mL

Therefore the required volume is 14.25 mL